# Attention as an Adaptive Filter

### v1

@@ 3.2 Complexity
The above tensor form of attention involves element-wise multiplication of $d\times m\times m$ tensors, resulting in time complexity $\mathcal{O}(m^{2}d+md^{2})$, which is the same as vanilla attention, but slower in practice due to less-optimized element-wise operations. The time complexity to propagate the precision matrices $\mathbf{P}_{ij}^{C}$ through the differential Lyapunov equation in the full (non-diagonalizable case) is $\mathcal{O}(d^{3})$, plus another $\mathcal{O}(d^{3})$ to invert each of these matrices, for each of the $\mathcal{O}(m^{2})$ time differences. Hence, the total time complexity is $\mathcal{O}(m^{2}d^{3})$. In contrast, for systems that admit a closed-form solution, the propagation and matrix inversions become $\mathcal{O}(d)$, so the total time complexity is $\mathcal{O}(m^{2}d)$. If the measurements are sampled at equal time intervals, the time complexity reduces to $\mathcal{O}(md)$.
@@ 3.3 Convolutional Representation
However, we must still store the full tensors of pulled-forward estimates $\mathcal{\hat{Z}}_{\mathbf{k}}[k,i,j]$ and $\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]$, which are of size $d_{k}\times m\times m$ and $d_{v}\times m\times m$, respectively.
@@ 3.5 Factorizing the Residuals
$\bm{\Phi}_{\mathbf{qk}}[k,i]=e^{\bm{\lambda}_{\mathbf{qk}}t_{i}},\quad\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]=e^{-\bm{\lambda}_{\mathbf{qk}}t_{j}}\mathbf{Z}_{\mathbf{k}}[k,j],\quad\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]=e^{-\bm{\lambda}_{\mathbf{qk}}t_{i}}\mathbf{Z}_{\mathbf{q}}[k,i],$
@@
We need to compute the robust weights:
@@
$\sum_{k}\|\bm{\Phi}_{\mathbf{qk}}[k,i]\|^{2}\left(\mathbf{\bar{\hat{Z}}}_{\mathbf{q}}[k,i]-\mathbf{\bar{\hat{Z}}}_{\mathbf{k}}[k,j]\right)\cdot\mathcal{P}^{C}_{\mathbf{qk}}[k,i,j]\cdot\left(\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]-\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)$
(where the overbar denotes element-wise complex conjugation). We can avoid constructing $\mathcal$$\hat${R}} by expanding the product, and $\mathcal{P}^{C}_{\mathbf{qk}}[k,i,j]$ by using the kernel $\mathcal{K}_{\mathbf{qk}}^{P}$:
$\mathbf{D}[i,j]=\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|]\|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\|^{2}+\sum_{k}\|\bm{\Phi}_{\mathbf{qk}}[k,i]\|^{2}\Big{[}\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|]\|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\|^{2}$
$-2\mathrm{Re}\left(\mathbf{\bar{\hat{Z}}}_{\mathbf{q}}[k,i]\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|]\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)\Big{]}$
But computing the product $\mathbf{\bar{\hat{Z}}}_{\mathbf{q}}[k,i]\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|]\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]$ still requires constructing a $d_{k}\times m\times m$ tensor.
@@ 3.6 Isotropic Decay and Noise
This corresponds to a system with state matrix $\mathbf{A}=\alpha\mathbf{I}+\bm{\Omega}$, where $\bm{\Omega}\in\mathbb{R}^{d\times d}$ is diagonalizable over $\mathbb{C}$, and has strictly imaginary eigenvalues, i.e. $\bm{\Omega}=\mathbf{S}\bm{\Lambda}_{\Omega}\mathbf{S}^{-1}$, where $\mathbf{S}\in\mathbb{C}^{d\times d}$, $\bm{\Lambda}_{\Omega}=\text{diag}(\lambda_{\Omega,1},\dots\lambda_{\Omega,d})$, with $\lambda_{\Omega,k}\in i\mathbb{R}$. We also assume that the noise is isotropic, i.e. that the noise parameters are scalar multiples of identity: $\bm{\Lambda}_{C}=\lambda_{C}\mathbf{I},\bm{\Lambda}_{Q}=\sigma^{2}\mathbf{I},\bm{\Lambda}_{R}=\eta^{2}\mathbf{I}$.
@@
$\mathcal{K}^{V}(\tau)=\sigma_{V}^{2}(\tau)\mathbf{I}_{d},\quad\text{where:}\quad\sigma_{V}^{2}(\tau)=\lambda_{C}^{2}\sigma^{2}\frac{1-e^{-2\alpha\Delta t\tau}}{-2\alpha}+\eta^{2}e^{-2\alpha\Delta t\tau}$
@@
We can then pull $\mathbf{P}^{C}_{\mathbf{k}}\in\mathbb{R}^{m\times m}$ outside the sum:
$\mathbf{D}[i,j]=\mathbf{P}^{C}_{\mathbf{k}}[i,j]\sum_{k}\|\bm{\Phi}_{\mathbf{qk}}[k,i]\|^{2}\Big{[}\|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\|^{2}+\|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\|^{2}-2\mathrm{Re}\left(\mathbf{\bar{\hat{Z}}}_{\mathbf{q}}[k,i]\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)\Big{]}$
@@
$\sum_{k}\|\bm{\Phi}_{\mathbf{qk}}[k,i]\|^{2}\|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\|^{2}=\sum_{k}\big{(}\|\bm{\Phi}_{\mathbf{qk}}\|^{2}\odot\|\mathbf{\hat{Z}}_{\mathbf{q}}\|^{2}\big{)}\in\mathbb{R}^{m},$
$\sum_{k}\|\bm{\Phi}_{\mathbf{qk}}[k,i]\|^{2}\|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\|^{2}=\big{(}\|\bm{\Phi}_{\mathbf{qk}}\|^{2}\big{)}^{\top}\|\mathbf{\hat{Z}}_{\mathbf{k}}\|^{2}\in\mathbb{R}^{m\times m},$
$\sum_{k}\|\bm{\Phi}_{\mathbf{qk}}[k,i]\|^{2}\mathbf{\bar{\hat{Z}}}_{\mathbf{q}}[k,i]\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]=\big{(}\|\bm{\Phi}_{\mathbf{qk}}\|^{2}\odot\mathbf{\hat{Z}}_{\mathbf{q}}\big{)}^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}\in\mathbb{C}^{m\times m}$
@@
$\|\mathbf{R}_{\mathbf{qk}}\|^{2}:=\mathbf{1}_{m}\big{(}\big{(}\|\bm{\Phi}_{\mathbf{qk}}\|^{2}\odot\|\mathbf{\hat{Z}}_{\mathbf{q}}\|^{2}\big{)}^{\top}\mathbf{1}_{d}\big{)}^{\top}+\big{(}\|\bm{\Phi}_{\mathbf{qk}}\|^{2}\big{)}^{\top}\|\mathbf{\hat{Z}}_{\mathbf{k}}\|^{2}-2\mathrm{Re}\Big{(}\big{(}\|\bm{\Phi}_{\mathbf{qk}}\|^{2}\odot\mathbf{\hat{Z}}_{\mathbf{q}}\big{)}^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}\Big{)}$
@@
$\|\mathbf{R}_{\mathbf{qk}}\|^{2}=e^{2\alpha t_{i}}\sum_{k}\Big{[}e^{-2\alpha t_{i}}\|\mathbf{Z_{q}}[k,i]\|^{2}+e^{-2\alpha t_{j}}\|\mathbf{Z_{k}}[k,j]\|^{2}-2\mathrm{Re}\left(e^{-\alpha t_{i}}e^{-\alpha t_{j}}\mathbf{\bar{\tilde{Z}}}_{\mathbf{q}}[k,i]\mathbf{\tilde{Z}}_{\mathbf{k}}[k,j]\right)\Big{]}$
$\sum_{k}\|\mathbf{Z_{q}}[k,i]\|^{2}+e^{2\alpha(t_{i}-t_{j})}\sum_{k}\|\mathbf{Z_{k}}[k,j]\|^{2}-2e^{\alpha(t_{i}-t_{j})}\sum_{k}\mathrm{Re}\left(\mathbf{\bar{\tilde{Z}}}_{\mathbf{q}}[k,i]\mathbf{\tilde{Z}}_{\mathbf{k}}[k,j]\right)$
Let $\mathbf{1}_{m}\in\mathbb{R}^{m\times 1}$ and $\mathbf{1}_{d_{k}}\in\mathbb{R}^{d_{k}\times 1}$ be column vectors of $1$s. In matrix form:
$\|\mathbf{R}_{\mathbf{qk}}\|^{2}=\mathbf{1}_{m}\big{(}\mathbf{1}_{d_{k}}^{\top}|\mathbf{Z_{q}}|^{2}\big{)}+\mathbf{E_{qk}}^{2}\odot\big{(}\mathbf{1}_{d_{k}}^{\top}|\mathbf{Z_{k}}|^{2}\big{)}^{\top}\mathbf{1}_{m}^{\top}-2\mathbf{E_{qk}}\odot\mathrm{Re}\big{(}\mathbf{\tilde{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\tilde{Z}}_{\mathbf{k}}\big{)}$
(where the modulus is taken element-wise). So we need three matrix products and two element-wise multiplications, as opposed to dot-product attention’s one matrix product.
We can then compute the un-normalized attention matrix:
$\mathbf{\tilde{A}}=\frac{\mathbf{P}^{C}_{\mathbf{v}}}{1+\mathbf{P}^{C}_{\mathbf{k}}\odot\|\mathbf{R}_{\mathbf{qk}}\|^{2}/\nu}$
(where division is again taken element-wise.) The above formulation reduces the required memory to $\mathcal{O}(m^{2})$.
Define a score matrix: $\mathbf{L}=\log(\mathbf{P_{v}}^{C})-\log\left(1+\mathbf{P_{k}}^{C}\odot\|\mathbf{R}_{\mathbf{qk}}\|^{2}/\nu\right)$,
and a causal mask $\mathbf{M}_{\text{causal}}$.
We can then express the row-normalization using softmax:
$\mathbf{A}[i,j]=e^{\mathbf{L}[i,j]\odot\mathbf{M}_{\text{causal}}}\big{/}\sum_{j}e^{\mathbf{L}[i,j]\odot\mathbf{M}_{\text{causal}}}$.
If the query-key and value dynamics parameters are shared, $\mathbf{P}^{C}_{\mathbf{k}}=\mathbf{P}^{C}_{\mathbf{v}}:=\mathbf{P}^{C}$, so:
$\mathbf{L}=-\log\left(\mathbf{V}^{C}+\|\mathbf{R}_{\mathbf{qk}}\|^{2}/\nu\right)$
@@
$\mathbf{\bar{Z}_{v}}^{\dagger}=(\bm{\tilde{\Phi}}_{\mathbf{v}}^{+})^{\dagger}\odot(\mathbf{\hat{A}}\mathbf{\tilde{Z}}_{\mathbf{v}}^{\dagger})$
@@
$\mathbf{Z}_{\mathbf{v}}^{\text{pred}}=(e^{\alpha\Delta t}e^{\bm{\lambda}_{\Omega}\Delta t})\odot\mathbf{\bar{Z}_{v}}$
This factorization, which separates scalar decay/growth (magnitude) from oscillatory components (direction), reduces the memory requirement of the algorithm from $\mathcal{O}(m^{2}d)$ to $\mathcal{O}(m^{2}+md)$, as in vanilla attention, while maintaining the $\mathcal{O}(m^{2}d)$ compute time. This comes at the price of losing the second-order information provided by the full tensors $\mathcal{P}^{C}_{\mathbf{qk}}$ and $\mathcal{P}^{C}_{\mathbf{v}}$. Whereas $\mathcal{P}^{C}_{\mathbf{qk}},\mathcal{P}^{C}_{\mathbf{v}}$ represent the confidence of the $k$-th feature’s projection from time $t_{j}$ to time $t_{i}$, $\mathbf{P}^{C}_{\mathbf{k}},\mathbf{P}^{C}_{\mathbf{v}}$ represent measurements of overall confidence between query and key.
The assumption that the real part of the eigenvalues is a shared scalar serves two purposes: (a) it allows the cross term to be represented as a matrix multiplication, which reduces the memory requirement, and (b) it improves the exploding/vanishing gradient problem for $\alpha<0$, because rather than computing $e^{\alpha t_{i}}$ and $e^{-\alpha t_{j}}($ separately, we compute $e^{\alpha(t_{i}-t_{j})}$, and $t_{i}>t_{j}$ for causal systems.
An alternative approach to simplify the tensor precision-weighted average into a tractable form with $\mathcal{O}(m^{2}+md)$ memory is to factorize the precision kernel into a product of factors depending separately on indices $i$ and $j$. We briefly explore this approach in the Appendix.

### v2

@@ 3.2 Complexity
The naive tensor implementation of attention above involves element-wise multiplication of $d\times m\times m$ tensors, resulting in time complexity $\mathcal{O}(m^{2}d+md^{2})$, which is the same as vanilla attention, but slower in practice due to less-optimized element-wise operations. The time complexity to propagate the precision matrices $\mathbf{P}_{ij}^{C}$ through the differential Lyapunov equation in the full (non-diagonalizable case) is $\mathcal{O}(d^{3})$, plus another $\mathcal{O}(d^{3})$ to invert each of these matrices, for each of the $\mathcal{O}(m^{2})$ time differences. Hence, the total time complexity is $\mathcal{O}(m^{2}d^{3})$. In contrast, for systems that admit a closed-form solution, the propagation and matrix inversions become $\mathcal{O}(d)$, so the total time complexity is $\mathcal{O}(m^{2}d)$. If the measurements are sampled at equal time intervals, the time complexity reduces to $\mathcal{O}(md)$.
@@ 3.3 Convolutional Representation
However, we must still store the full tensors of pulled-forward estimates $\mathcal{\hat{Z}}_{\mathbf{k}}[k,i,j]$ and $\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]$, which are of size $d_{qk}\times m\times m$ and $d_{v}\times m\times m$, respectively.
@@ 3.5 Factorizing the Residuals
$\bm{\Phi}_{\mathbf{qk}}[k,i]:=e^{\bm{\lambda}_{\mathbf{qk}}t_{i}},\quad\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]:=e^{-\bm{\lambda}_{\mathbf{qk}}t_{i}}\mathbf{Z}_{\mathbf{q}}[k,i],\quad\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]:=e^{-\bm{\lambda}_{\mathbf{qk}}t_{j}}\mathbf{Z}_{\mathbf{k}}[k,j],$
@@
We need to compute the matrix of Mahalanobis distances:
$\mathbf{D}[i,j]:=\sum_{k}\mathcal{R}^{*}_{\mathbf{qk}}[k,i,j]\cdot\mathcal{P}^{C}_{\mathbf{qk}}[k,i,j]\cdot\mathcal{R}_{\mathbf{qk}}[k,i,j]$
@@
$\sum_{k}\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\left(\mathbf{\hat{Z}}^{*}_{\mathbf{q}}[k,i]-\mathbf{\hat{Z}}^{*}_{\mathbf{k}}[k,j]\right)\cdot\mathcal{P}^{C}_{\mathbf{qk}}[k,i,j]\cdot\left(\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]-\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)$
@@
$\mathbf{D}[i,j]=\sum_{k}\mathcal{K}^{C}_{\mathbf{qk}}[k,|i-j|]\cdot\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\cdot\Big[\big|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\big|^{2}+\big|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big|^{2}-2\mathrm{Re}\big(\mathbf{\hat{Z}}^{*}_{\mathbf{q}}[k,i]\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big)\Big]$
But computing the product $\mathbf{\hat{Z}}^{*}_{\mathbf{q}}[k,i]\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|]\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]$ still requires constructing a $d_{qk}\times m\times m$ tensor.
@@ 3.6 Isotropic Decay and Noise
This corresponds to a system with an isotropic plus skew-symmetric state matrix, $\mathbf{A}=\alpha\mathbf{I}+\bm{\Omega}$, where $\bm{\Omega}\in\mathbb{R}^{d\times d}$ is skew-symmetric ($\bm{\Omega}^{\top}=-\bm{\Omega}$) and diagonalizable over $\mathbb{C}$, that is $\bm{\Omega}=\mathbf{S}\bm{\Lambda}_{\Omega}\mathbf{S}^{-1}$, where $\mathbf{S}\in\mathbb{C}^{d\times d}$, $\bm{\Lambda}_{\Omega}=\text{diag}(\omega_{1},\dots\omega_{d})$, with $\omega_{k}\in i\mathbb{R}$. We also assume that the noise is isotropic, i.e. that the noise parameters are scalar multiples of identity: $\bm{\Lambda}_{C}=\lambda_{C}\mathbf{I},\bm{\Lambda}_{Q}=\sigma^{2}\mathbf{I},\bm{\Lambda}_{R}=\eta^{2}\mathbf{I}$.
@@
$\mathcal{K}^{V}(\tau)=\sigma_{V}^{2}(\tau)\mathbf{I}_{d},\quad\text{where:}\quad\sigma_{V}^{2}(\tau)=\big|\lambda_{C}\big|^{2}\sigma^{2}\frac{1-e^{-2\alpha\Delta t\tau}}{-2\alpha}+\eta^{2}e^{-2\alpha\Delta t\tau}$
@@
We can then pull $\mathbf{P}^{C}_{\mathbf{qk}}\in\mathbb{R}^{m\times m}$ outside the sum:
$\mathbf{D}[i,j]=\mathbf{P}^{C}_{\mathbf{qk}}[i,j]\sum_{k}\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\Big[\big|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\big|^{2}+\big|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big|^{2}-2\mathrm{Re}\left(\mathbf{\hat{Z}}^{*}_{\mathbf{q}}[k,i]\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)\Big]$
@@
$\sum_{k}\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\big|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\big|^{2}=\sum_{k}\big(\big|\bm{\Phi}_{\mathbf{qk}}\big|^{2}\odot\big|\mathbf{\hat{Z}}_{\mathbf{q}}\big|^{2}\big)\in\mathbb{R}^{m\times 1},$
$\sum_{k}\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\big|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big|^{2}=\big(\big|\bm{\Phi}_{\mathbf{qk}}\big|^{2}\big)^{\top}\big|\mathbf{\hat{Z}}_{\mathbf{k}}\big|^{2}\in\mathbb{R}^{m\times m},$
$\sum_{k}\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\mathbf{\hat{Z}}^{*}_{\mathbf{q}}[k,i]\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]=\big(\big|\bm{\Phi}_{\mathbf{qk}}\big|^{2}\odot\mathbf{\hat{Z}}_{\mathbf{q}}\big)^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}\in\mathbb{C}^{m\times m}$
@@
$\big|\mathbf{R}_{\mathbf{qk}}\big|^{2}:=\mathbf{1}_{m}\big(\big(\big|\bm{\Phi}_{\mathbf{qk}}\odot\mathbf{\hat{Z}}_{\mathbf{q}}\big|^{2}\big)^{\top}\mathbf{1}_{d}\big)^{\top}+\big(\big|\bm{\Phi}_{\mathbf{qk}}^{\top}\odot\mathbf{\hat{Z}}_{\mathbf{k}}\big|^{2}-2\mathrm{Re}\Big(\big(\bm{\Phi}_{\mathbf{qk}}\odot\mathbf{\hat{Z}}_{\mathbf{q}}\big)^{\dagger}\big(\bm{\Phi}_{\mathbf{qk}}\odot\mathbf{\hat{Z}}_{\mathbf{k}}\big)\Big)$
@@
$\big|\mathbf{R}_{\mathbf{qk}}[i,j]\big|^{2}=e^{2\alpha t_{i}}\sum_{k}\Big[e^{-2\alpha t_{i}}\big|\mathbf{Z_{q}}[k,i]\big|^{2}+e^{-2\alpha t_{j}}\big|\mathbf{Z_{k}}[k,j]\big|^{2}-2\mathrm{Re}\left(e^{-\alpha t_{i}}e^{-\alpha t_{j}}\mathbf{\tilde{Z}}^{*}_{\mathbf{q}}[k,i]\mathbf{\tilde{Z}}_{\mathbf{k}}[k,j]\right)\Big]$
$\sum_{k}\big|\mathbf{Z_{q}}[k,i]\big|^{2}+e^{2\alpha(t_{i}-t_{j})}\sum_{k}\big|\mathbf{Z_{k}}[k,j]\big|^{2}-2e^{\alpha(t_{i}-t_{j})}\sum_{k}\mathrm{Re}\left(\mathbf{\tilde{Z}}^{*}_{\mathbf{q}}[k,i]\mathbf{\tilde{Z}}_{\mathbf{k}}[k,j]\right)$
Let $\mathbf{1}_{m}\in\mathbb{R}^{m\times 1}$ and $\mathbf{1}_{d_{qk}}\in\mathbb{R}^{d_{qk}\times 1}$ be column vectors of $1$s. In matrix form:
$\big|\mathbf{R}_{\mathbf{qk}}\big|^{2}=\mathbf{1}_{m}\big(\mathbf{1}_{d_{qk}}^{\top}|\mathbf{Z_{q}}|^{2}\big)+\mathbf{E_{qk}}^{2}\odot\big(\mathbf{1}_{d_{qk}}^{\top}|\mathbf{Z_{k}}|^{2}\big)^{\top}\mathbf{1}_{m}^{\top}-2\mathbf{E_{qk}}\odot\mathrm{Re}\big(\mathbf{\tilde{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\tilde{Z}}_{\mathbf{k}}\big)$
(where the modulus is taken element-wise).
Defining a score matrix $\mathbf{L}=\log(\mathbf{P}_{\mathbf{v}}^{C})-\log\left(\mathbf{I}+\mathbf{P}^{C}_{\mathbf{qk}}\odot\big|\mathbf{R}_{\mathbf{qk}}\big|^{2}/\nu\right)$, and a causal mask $\mathbf{M}_{\text{causal}}$, we can then express the row-normalization using softmax:
$\mathbf{A}[i,j]=e^{\mathbf{L}[i,j]\odot\mathbf{M}_{\text{causal}}}\big/\sum_{j}e^{\mathbf{L}[i,j]\odot\mathbf{M}_{\text{causal}}}$
@@
Whereas for the query-key dynamics, the isotropic decay condition $\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})=\alpha\mathbf{1}$ is necessary to reduce the memory requirement, there is no such constraint on $\bm{\lambda}_{\mathbf{v}}$. However, we face another problem: for $\mathrm{Re(}\bm{\lambda}_{\mathbf{v}})\leq\mathbf{0}$, $\mathbf{\hat{Z}_{v}}$ grows exponentially and $\bm{\Phi}_{\mathbf{v}}$ decays exponentially. However, if we again assume isotropic decay $\mathrm{Re}(\bm{\lambda}_{\mathbf{v}})=\alpha\mathbf{1}$, we can re-factorize $\mathbf{\bar{Z}_{v}}[k,i]$ such that this problem is avoided.
@@
$\mathbf{\bar{Z}}_{\mathbf{v}}^{\dagger}=(\bm{\tilde{\Phi}}_{\mathbf{v}}^{+})^{\dagger}\odot(\mathbf{\hat{A}}\mathbf{\tilde{Z}}_{\mathbf{v}}^{\dagger})$
@@
$\mathbf{Z}_{\mathbf{v}}^{+}=e^{\bm{\lambda}\Delta t}\odot\mathbf{\bar{Z}_{v}}$
This factorization, which separates scalar decay/growth (magnitude) from oscillatory components (direction), reduces the memory requirement of the algorithm from $\mathcal{O}(m^{2}d)$ to $\mathcal{O}(m^{2}+md)$, as in vanilla attention, while maintaining the $\mathcal{O}(m^{2}d)$ compute time. This comes at the price of losing the second-order information provided by the full tensors $\mathcal{P}^{C}_{\mathbf{qk}}$ and $\mathcal{P}^{C}_{\mathbf{v}}$. Whereas $\mathcal{P}^{C}_{\mathbf{qk}},\mathcal{P}^{C}_{\mathbf{v}}$ represent the confidence of the $k$-th feature’s projection from time $t_{j}$ to time $t_{i}$, $\mathbf{P}^{C}_{\mathbf{qk}},\mathbf{P}^{C}_{\mathbf{v}}$ represent measurements of overall confidence between query and key.
Computing the residual matrix $\big|\mathbf{R}_{\mathbf{qk}}\big|^{2}$ requires additional vector summations, broadcasting, and element-wise multiplications, as opposed to dot-product attention’s single matrix product. However, computing $|\mathbf{Z_{q}|}$ and $|\mathbf{Z_{k}|}$ is only $\mathcal{O}(d_{qk}m)$, so it is negligible overhead compared to the $\mathcal{O}(d_{qk}m^{2})$ dot product computation.
The assumption that the real part of the eigenvalues is a shared scalar serves two purposes: (a) it allows the cross term to be represented as a matrix multiplication, which reduces the memory requirement, and (b) it improves the exploding/vanishing gradient problem for $\alpha<0$, because rather than computing $e^{\alpha t_{i}}$ and $e^{-\alpha t_{j}}$ separately, we compute $e^{\alpha(t_{i}-t_{j})}$, which is stable, since $t_{i}>t_{j}$ for causal systems.
[…]
Practically, the differences are:
  • AFA uses complex-valued $\mathbf{W_{q}}$, $\mathbf{W_{k}}$, $\mathbf{W_{v}}$, $\mathbf{W_{p}}$, whereas RetNet uses real-valued projections.
  • AFA also applies the matrix exponential to the value projections, whereas RetNet does not, keeping its values static.
  • AFA incorporates uncertainty priors into attention using propagated covariance kernels.
3.7 Normalized Queries & Keys
If we assume normalized queries and keys, $\sum_{k}\big|\mathbf{Z_{q}}[k,i]\big|^{2}=1$, $\sum_{k}\big|\mathbf{Z_{k}}[k,i]\big|^{2}=1$, the expression for the squared residual simplifies:
$\big|\mathbf{R}_{\mathbf{qk}}[i,j]\big|^{2}=1+e^{2\alpha(t_{i}-t_{j})}-2e^{\alpha(t_{i}-t_{j})}\sum_{k}\mathrm{Re}\left(\mathbf{\tilde{Z}}^{*}_{\mathbf{q}}[k,i]\mathbf{\tilde{Z}}_{\mathbf{k}}[k,j]\right)$
Or in matrix, form:
$\big|\mathbf{R}_{\mathbf{qk}}\big|^{2}=\mathbf{I}+\mathbf{E_{qk}}^{2}-2\mathbf{E_{qk}}\odot\mathrm{Re}\big(\mathbf{\tilde{Z}}_{\mathbf{q}}^{\dagger}\mathbf{\tilde{Z}}_{\mathbf{k}}\big)$
If the query-key and value dynamics parameters are shared, $\mathbf{P}^{C}_{\mathbf{qk}}=\mathbf{P}^{C}_{\mathbf{v}}:=\mathbf{P}^{C}$, so:
$\mathbf{\tilde{A}}=\frac{\nu\mathbf{I}}{\nu\mathbf{V}^{C}+\big|\mathbf{R}_{\mathbf{qk}}\big|^{2}}$
With both of these simplifications, the term in the denominator is:
$(\nu\mathbf{V}^{C}+\big|\mathbf{R}_{\mathbf{qk}}\big|^{2})[i,j]=\nu\big|\lambda_{C}\big|^{2}\sigma^{2}\frac{1-e^{-2\alpha\Delta t_{ij}}}{-2\alpha}+\eta^{2}e^{-2\alpha\Delta t_{ij}}+1+e^{-2\alpha\Delta t_{ij}}$
$-2e^{-\alpha\Delta t_{ij}}\sum_{k}\mathrm{Re}\left(\mathbf{\tilde{Z}}^{*}_{\mathbf{q}}[k,i]\mathbf{\tilde{Z}}_{\mathbf{k}}[k,j]\right)$
Collecting terms and letting $a=1+\frac{\nu|\lambda_{C}|^{2}\sigma^{2}}{-2\alpha}$, $b=1+\eta^{2}-\frac{\nu|\lambda_{C}|^{2}\sigma^{2}}{-2\alpha}$, this can be written as:
$(\nu\mathbf{V}^{C}+\big|\mathbf{R}_{\mathbf{qk}}\big|^{2})[i,j]=a+be^{-2\alpha\Delta t_{ij}}-2e^{-\alpha\Delta t_{ij}}\sum_{k}\mathrm{Re}\left(\mathbf{\tilde{Z}}^{*}_{\mathbf{q}}[k,i]\mathbf{\tilde{Z}}_{\mathbf{k}}[k,j]\right)$
The attention score denominator is now explicitly decomposed into a content-independent positional decay term and a content-dependent similarity term modulated by the time difference.